# Improving Robustness to Model Inversion Attacks via Sparse Coding Architectures

### v1

@@ Appendix 0.D Additional baseline tuning
[…]
Accuracy
MNIST
Titcombe et al.[71]-$1.0$
$22.63$
$0.503$
$66.40$
$0.980$
SCA0.1
$\mathbf{16.17}$
$\mathbf{0.109}$
$\mathbf{227.4}$
$\mathbf{0.988}$
SCA0.25
$\mathbf{17.40}$
$\mathbf{0.058}$
$\mathbf{301.6}$
$\mathbf{0.980}$
SCA0.5
$\mathbf{14.98}$
$\mathbf{0.044}$
$\mathbf{307.7}$
$\mathbf{0.975}$
Fashion
Titcombe et al.[71]-$1.0$
$18.36$
$0.408$
$80.80$
$0.878$
MNIST
SCA0.1
$\mathbf{18.11}$
$\mathbf{0.154}$
$\mathbf{171.1}$
$\mathbf{0.904}$
SCA0.25
$\mathbf{17.74}$
$\mathbf{0.188}$
$\mathbf{203.8}$
$\mathbf{0.896}$
SCA0.5
$\mathbf{17.15}$
$\mathbf{0.134}$
$\mathbf{270.4}$
$\mathbf{0.879}$
CIAFR10
Titcombe et al.[71]-$1.0$
$14.27$
$0.259$
$171.6$
$0.786$
SCA0.1
$\mathbf{13.18}$
$\mathbf{0.002}$
$\mathbf{174.2}$
$\mathbf{0.758}$
SCA0.25
$\mathbf{13.07}$
$\mathbf{0.002}$
$\mathbf{181.2}$
$\mathbf{0.742}$
SCA0.5
$\mathbf{12.88}$
$\mathbf{0.002}$
$\mathbf{375.3}$
$\mathbf{0.739}$
@@ Appendix 0.E Stability analysis of SCA
[…]
Accuracy
CelebA
No-Defense
$11.42\pm 2.44$
$0.613\pm 0.29$
$292.9\pm 81.5$
$0.721\pm 0.04$
Gaussian-Noise
$10.87\pm 2.25$
$0.614\pm 0.30$
$296.5\pm 73.3$
$0.624\pm 0.03$
GAN
$11.02\pm 1.82$
$0.600\pm 0.29$
$301.4\pm 92.4$
$0.613\pm 0.02$
Gong et al.[22]++
$10.84\pm 1.94$
$0.556\pm 0.28$
$301.0\pm 81.4$
$0.658\pm 0.02$
Titcombe et al.[71]
$10.76\pm 2.37$
$0.557\pm 0.24$
$345.5\pm 86.1$
$0.643\pm 0.01$
Gong et al.[22]
$10.91\pm 1.88$
$0.560\pm 0.29$
$304.5\pm 82.5$
$0.616\pm 0.01$
Peng et al.[58]
$10.17\pm 2.32$
$0.491\pm 0.24$
$399.1\pm 55.3$
$0.667\pm 0.04$
Hayes et al.[25]
$10.16\pm 1.95$
$0.535\pm 0.25$
$320.8\pm 79.0$
$0.601\pm 0.02$
Wang et al.[78]
$10.39\pm 2.55$
$0.505\pm 0.24$
$341.7\pm 74.2$
$0.669\pm 0.05$
Sparse-Std
$9.78\pm 2.13$
$0.485\pm 0.24$
$367.3\pm 44.7$
$0.663\pm 0.03$
SCA0.1
$\mathbf{9.56}\pm 2.30$
$\mathbf{0.454}\pm 0.25$
$\mathbf{396.6}\pm 45.0$
$\mathbf{0.659}\pm 0.04$
SCA0.25
$\mathbf{9.27}\pm 2.06$
$\mathbf{0.452}\pm 0.25$
$\mathbf{412.8}\pm 53.7$
$\mathbf{0.661}\pm 0.05$
SCA0.5
$\mathbf{9.12}\pm 2.68$
$\mathbf{0.368}\pm 0.24$
$\mathbf{421.7}\pm 49.9$
$\mathbf{0.653}\pm 0.04$
Dataset
Defense
[…]
Accuracy
Medical
No-Defense
$30.17\pm 0.90$
$0.912\pm 0.01$
$12.40\pm 8.69$
$0.998\pm 0.01$
MNIST
Gaussian-Noise
$27.00\pm 1.30$
$0.828\pm 0.05$
$17.29\pm 11.9$
$0.886\pm 0.06$
GAN
$25.05\pm 2.78$
$0.699\pm 0.03$
$29.08\pm 20.5$
$0.995\pm 0.01$
Gong et al.[22]++
$20.37\pm 1.65$
$0.451\pm 0.03$
$44.68\pm 30.9$
$0.871\pm 0.01$
Titcombe et al.[71]
$20.51\pm 0.28$
$0.574\pm 0.01$
$28.23\pm 1.65$
$0.805\pm 0.06$
Gong et al.[22]
$23.65\pm 1.07$
$0.605\pm 0.09$
$37.16\pm 26.2$
$0.757\pm 0.03$
Peng et al.[58]
$17.42\pm 2.87$
$0.519\pm 0.22$
$65.39\pm 32.8$
$0.866\pm 0.08$
Hayes et al.[25]
$19.57\pm 1.08$
$0.003\pm 0.01$
$155.0\pm 92.5$
$0.847\pm 0.08$
Wang et al.[78]
$17.89\pm 2.09$
$0.463\pm 0.08$
$101.8\pm 66.5$
$0.829\pm 0.08$
Sparse-Std
$13.49\pm 0.29$
$0.158\pm 0.09$
$203.4\pm 92.2$
$0.865\pm 0.05$
SCA0.1
$\mathbf{12.46}\pm 0.30$
$\mathbf{0.006}\pm 0.01$
$\mathbf{231.8}\pm 124$
$\mathbf{0.858}\pm 0.08$
SCA0.25
$\mathbf{11.89}\pm 0.35$
$\mathbf{0.008}\pm 0.01$
$\mathbf{254.1}\pm 153$
$\mathbf{0.850}\pm 0.08$
SCA0.5
$\mathbf{11.19}\pm 0.11$
$\mathbf{0.001}\pm 0.01$
$\mathbf{276.9}\pm 97.0$
$\mathbf{0.841}\pm 0.08$
@@ Appendix 0.F Compute time
Appendix 0.G Robustness of sparse coding layers: UMap

### v2

@@ Appendix 0.D Additional baseline tuning
MNIST Titcombe et al.[75]-$1.0$ $22.63$ $0.503$ $66.40$ $0.980$
SCA0.1 $\mathbf{16.17}$ $\mathbf{0.109}$ $\mathbf{227.4}$ $\mathbf{0.988}$
SCA0.25 $\mathbf{17.40}$ $\mathbf{0.058}$ $\mathbf{301.6}$ $\mathbf{0.980}$
SCA0.5 $\mathbf{14.98}$ $\mathbf{0.044}$ $\mathbf{307.7}$ $\mathbf{0.975}$
Fashion Titcombe et al.[75]-$1.0$ $18.36$ $0.408$ $80.80$ $0.878$
MNIST SCA0.1 $\mathbf{18.11}$ $\mathbf{0.154}$ $\mathbf{171.1}$ $\mathbf{0.904}$
SCA0.25 $\mathbf{17.74}$ $\mathbf{0.188}$ $\mathbf{203.8}$ $\mathbf{0.896}$
SCA0.5 $\mathbf{17.15}$ $\mathbf{0.134}$ $\mathbf{270.4}$ $\mathbf{0.879}$
CIAFR10 Titcombe et al.[75]-$1.0$ $14.27$ $0.259$ $171.6$ $0.786$
SCA0.1 $\mathbf{13.18}$ $\mathbf{0.002}$ $\mathbf{174.2}$ $\mathbf{0.758}$
SCA0.25 $\mathbf{13.07}$ $\mathbf{0.002}$ $\mathbf{181.2}$ $\mathbf{0.742}$
SCA0.5 $\mathbf{12.88}$ $\mathbf{0.002}$ $\mathbf{375.3}$ $\mathbf{0.739}$
@@ Appendix 0.E Stability analysis of SCA
CelebA No-Defense $11.42\pm 2.44$ $0.613\pm 0.29$ $292.9\pm 81.5$ $0.721\pm 0.04$
Gaussian-Noise $10.87\pm 2.25$ $0.614\pm 0.30$ $296.5\pm 73.3$ $0.624\pm 0.03$
GAN $11.02\pm 1.82$ $0.600\pm 0.29$ $301.4\pm 92.4$ $0.613\pm 0.02$
Gong et al.[25]++ $10.84\pm 1.94$ $0.556\pm 0.28$ $301.0\pm 81.4$ $0.658\pm 0.02$
Titcombe et al.[75] $10.76\pm 2.37$ $0.557\pm 0.24$ $345.5\pm 86.1$ $0.643\pm 0.01$
Gong et al.[25] $10.91\pm 1.88$ $0.560\pm 0.29$ $304.5\pm 82.5$ $0.616\pm 0.01$
Peng et al.[62] $10.17\pm 2.32$ $0.491\pm 0.24$ $399.1\pm 55.3$ $0.667\pm 0.04$
Hayes et al.[28] $10.16\pm 1.95$ $0.535\pm 0.25$ $320.8\pm 79.0$ $0.601\pm 0.02$
Wang et al.[82] $10.39\pm 2.55$ $0.505\pm 0.24$ $341.7\pm 74.2$ $0.669\pm 0.05$
Sparse-Std $9.78\pm 2.13$ $0.485\pm 0.24$ $367.3\pm 44.7$ $0.663\pm 0.03$
SCA0.1 $\mathbf{9.56}\pm 2.30$ $\mathbf{0.454}\pm 0.25$ $\mathbf{396.6}\pm 45.0$ $\mathbf{0.659}\pm 0.04$
SCA0.25 $\mathbf{9.27}\pm 2.06$ $\mathbf{0.452}\pm 0.25$ $\mathbf{412.8}\pm 53.7$ $\mathbf{0.661}\pm 0.05$
SCA0.5 $\mathbf{9.12}\pm 2.68$ $\mathbf{0.368}\pm 0.24$ $\mathbf{421.7}\pm 49.9$ $\mathbf{0.653}\pm 0.04$
@@
Medical No-Defense $30.17\pm 0.90$ $0.912\pm 0.01$ $12.40\pm 8.69$ $0.998\pm 0.01$
MNIST Gaussian-Noise $27.00\pm 1.30$ $0.828\pm 0.05$ $17.29\pm 11.9$ $0.886\pm 0.06$
GAN $25.05\pm 2.78$ $0.699\pm 0.03$ $29.08\pm 20.5$ $0.995\pm 0.01$
Gong et al.[25]++ $20.37\pm 1.65$ $0.451\pm 0.03$ $44.68\pm 30.9$ $0.871\pm 0.01$
Titcombe et al.[75] $20.51\pm 0.28$ $0.574\pm 0.01$ $28.23\pm 1.65$ $0.805\pm 0.06$
Gong et al.[25] $23.65\pm 1.07$ $0.605\pm 0.09$ $37.16\pm 26.2$ $0.757\pm 0.03$
Peng et al.[62] $17.42\pm 2.87$ $0.519\pm 0.22$ $65.39\pm 32.8$ $0.866\pm 0.08$
Hayes et al.[28] $19.57\pm 1.08$ $0.003\pm 0.01$ $155.0\pm 92.5$ $0.847\pm 0.08$
Wang et al.[82] $17.89\pm 2.09$ $0.463\pm 0.08$ $101.8\pm 66.5$ $0.829\pm 0.08$
Sparse-Std $13.49\pm 0.29$ $0.158\pm 0.09$ $203.4\pm 92.2$ $0.865\pm 0.05$
SCA0.1 $\mathbf{12.46}\pm 0.30$ $\mathbf{0.006}\pm 0.01$ $\mathbf{231.8}\pm 124$ $\mathbf{0.858}\pm 0.08$
SCA0.25 $\mathbf{11.89}\pm 0.35$ $\mathbf{0.008}\pm 0.01$ $\mathbf{254.1}\pm 153$ $\mathbf{0.850}\pm 0.08$
SCA0.5 $\mathbf{11.19}\pm 0.11$ $\mathbf{0.001}\pm 0.01$ $\mathbf{276.9}\pm 97.0$ $\mathbf{0.841}\pm 0.08$
@@ Appendix 0.F Compute time
Appendix 0.G Ablations: Tuning SCA
Observe that our SCA outperforms SOTA defense baselines in robustness even without any tuning of parameters. However, tuning the hyper-parameters can boost the accuracy further, e.g., we use kernel size as default 5 for all experiments. Increasing the kernel from 5 to 7 can improve SCA accuracies beyond. While our goal is to capture the essence of the SCA itself in terms of robustness, we explore a little bit on further possible improvements on accuracy scores. We consider the lowest robust SCA, i.e., SCA0.1 for the tuning of kernel size, and we present the comparisons of accuracies between SCA0.1 and Tuned SCA0.1 in Table 15.
Table 15: Comparison of Accuracy Scores among our unoptimized SCA0.1 and Tuned SCA0.1 (kernel: 5$\rightarrow$ 7) in all 3 setups on CelebA and Medical MNIST datasets.
[…]
CelebA plug and play $0.726$ $\mathbf{0.730}$
end to end $0.748$ $\mathbf{0.751}$
split $0.745$ $\mathbf{0.759}$
Medical plug and play $0.888$ $\mathbf{0.899}$
MNIST end to end $0.888$ $\mathbf{0.996}$
split $0.946$ $\mathbf{0.967}$
@@ Appendix 0.G Ablations: Tuning SCA
Appendix 0.H Robustness of sparse coding layers: UMap